### v2

@@ 4.1 Causal clustering: algorithm
Equation (14) solves a weighted min-cut problem with an additional penalization term that depends on the variance of the clusters’ size. The optimization problem in Equation (14) is similar in spirit to, though different from, the minimum normalized-cut problem (e.g., Shi and Malik, 2000; Ling and Strohmer, 2020), whose objective is $\sum_{k=1}^{K}\sum_{i:c(i)=k}|j\in\mathcal{N}_{i}:c(i)\neq c(j)|/\sum_{i:c(i)=k}|\mathcal{N}_{i}|.$ Different from our proposal, standard min-cut problems use a different denominator (since the objective is not motivated by the bias of treatment effect estimators), and do not account for the additional variance component. Here, researchers should balance the bias and variance in the choice of the clusters. To the best of our knowledge, the problem in Equation (14) has not been studied before.
@@
In the following lines we exploit Corollary 4.1 to find the clustering that minimizes Equation (14) over a large space of clusters. Let $\mathbf{V}=\text{diag}(\mathbf{A}\mathbf{1}_{n})=\text{diag}(|\mathcal{N}_{1}(1)|,\ldots,|\mathcal{N}_{n}(1)|)$. Define the left-normalized Laplacian
@@
Let tr()normal-$\mathrm{tr}(\cdot)$ denote the trace operator. $c^{\star}$ minimizes Equation (14) if and only if
$c^{\star}\in\mathrm{arg}\max_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}\mathrm{tr}\Big{(}(n\mathbf{L}-\xi_{n}\mathbf{1}_{n}\mathbf{1}_{n}^{T})\mathbf{M}_{c}(K)\mathbf{M}_{c}^{T}(K)\Big{)}.$
@@
See Appendix C.9. ∎
Theorem 4.1 formalizes the optimization problem as a trace-optimization program for a fixed number of clusters $K$. Theorem 4.1 does not characterize a convex optimization program, but it provides a natural starting point to study convex relaxations of the proposed optimization problem. To obtain a convex relaxation, we relax the constraint on the matrix $\mathbf{X}(K)=\mathbf{M}_{c}(K)\mathbf{M}_{c}^{T}(K)$. We propose solving the following semi-definite programming (SDP) problem (for a given number of clusters $K$):
$\max_{\mathbf{X}(K)}\text{tr}(\mathbf{L}_{\xi_{n}}\mathbf{X}(K)),\quad\text{s.t.}\quad\text{diag}(\mathbf{X}(K))=\mathbf{1}_{n},\,\,\mathbf{X}(K)\succeq 0,\quad\mathbf{L}_{\xi_{n}}=n\mathbf{L}-\xi_{n}\mathbf{1}_{n}\mathbf{1}_{n}^{T}.$ (16)
where Equation (16) defines a sequence of semi-definite optimization programs, each indexed by the number of clusters $K$. The main distinction between Equation (14), and Equation (16) is that the matrix $\mathbf{X}(K)$ must be positive-definite instead of a matrix with binary entries. Let $\hat{\mathbf{X}}(K)$ be the solution of (16), for given $K$, that can be obtained using off-the-shelf optimization routines. We then apply the $K$-means algorithm on the first $K$ eigenvectors of $\hat{\mathbf{X}}(K)$ to retrieve the mapping $c$, whose properties are well studied in the literature (Von Luxburg, 2007). Finally, we compare solutions for different values of $K$ and choose the clustering with the largest objective.
In summary, the complete algorithm (Algorithm 1) solves a sequence of semi-definite trace-optimization problems, each indexed by a different value of $K$, and reports the clustering (and corresponding number of clusters) that leads to the largest objective.
@@
Remark 8 (Spectral relaxation).
@@
$\max_{\mathbf{U}}\text{tr}(\mathbf{L}_{\xi_{n}}\mathbf{U}(K)\mathbf{U}^{T}(K)),\quad\text{s.t.}\,\,\mathbf{U}^{T}(K)\mathbf{U}(K)=\mathbf{I}_{K},$
@@
$\text{tr}(\mathbf{L}_{\xi_{n}}\mathbf{U}(K)\mathbf{U}^{T}(K))=\text{tr}\left(\frac{\mathbf{L}_{\xi_{n}}+\mathbf{L}_{\xi_{n}}^{T}}{2}\mathbf{U}\mathbf{U}^{T}\right)$
The solution to the above problem $\hat{\mathbf{U}}(K)\in\mathbb{R}^{n\times K}$ is given by the matrix of top-$K$ eigenvectors of $(\mathbf{L}_{\xi_{n}}+\mathbf{L}_{\xi_{n}}^{T})/2$. Then we can perform a constrained $K$-means algorithm with equal-sized clusters to recover the clusters (Bradley et al., 2000). ∎
Algorithm 1 Causal Clustering
1:Adjacency matrix $\mathbf{A}$, $\underline{K}$, $\bar{K}$ smallest and largest number of clusters, $\xi_{n}$.
2:for $K\in\{\underline{K},\cdots,\bar{K}\}$ do
3:a: Solve Equation (16) and obtain $\hat{\mathbf{X}}(K)$ as the minimized of Equation (16)
4:b: Retrive the clusters $c_{K}$ via $K$-means algorithm on the first $K$ eigenvectors of $\hat{\mathbf{X}}(K)$
5:c: Compute the objective function corresponding to the chosen clustering in (14)
6:end for
7:return Clustering with the lowest objective function in (14).
@@ 5 Empirical illustration and numerical studies
In this section, we illustrate the properties of the procedure in two empirical applications, one using unique data from the Facebook friendship and messaging graphs and one using data from an experiment conducted in rural China by Cai et al. (2015). We provide further evidence supporting our theoretical analysis using simulated networks in Section 5.3.
5.1 Clustering on Facebook graphs
We evaluate our procedure on the universe of users using Facebook and Messenger.
5.1.1 Clustering algorithms, bias and variance
We consider two clustering algorithms implemented at scale on social networks owned by Meta: Louvain algorithm (Blondel et al., 2008), and Balanced Partitioning (Kabiljo et al., 2017). Louvain and Balanced Partitioning clustering produces a hierarchical clustering structure with a growing number of clusters. We consider three “types” of such algorithms, corresponding to three levels in the clustering structure hierarchy (i.e., different numbers of clusters), defined as Type 1, 2, and 3, respectively. For Louvain, higher-order types denote more clusters, whereas it is the opposite for Balanced Partitioning. For each clustering, we also report $\log(n/K_{n})$, the log ratio of the number of clusters over the population size.
@@ 5.1.1 Clustering algorithms, bias and variance
We report the bias and variance, with the variance weighted by a parameter $\xi_{n}$, defined respectively as
$\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}\mathcal{N}_{i}\bigcap\Big{\{}j:c(i)\neq c(i)\Big{\}}\Big{|},\quad\frac{\xi_{n}}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}.$
5.1.2 Comparisons between different algorithms
@@
Table 3 collects each of these algorithms’ worst-case bias and variance and the number of clusters per individual. A larger number of clusters increases the bias but decreases the variance throughout all graphs. (The bias increases for sparser networks because the denominator $|\mathcal{N}_{i}|$ decreases.) Louvain algorithm dominates Balanced Partitioning uniformly in variance and bias, suggesting that practitioners may prefer Louvain over Balanced Partitioning.
@@
We study different types of Louvain algorithms (each having a different number of clusters) in Figure 2. Figure 2 reports the worst-case mean squared error for different graphs, Louvain algorithms, and degrees of sparsity. For $\xi_{n}\approx 1$, the Louvain algorithm with the largest number of clusters (Type 3) dominates the other two algorithms in all but one case.
@@
As $\xi_{n}$ increases, Louvain clustering with the largest number of clusters (Type 3) is optimal for a dense graph but sub-optimal as the graph becomes more sparse. This result is intuitive: for denser networks, a larger number of clusters may best control the estimator’s bias relative to the bias induced by other clustering algorithms. However, we observe that Type 1 is the most robust in terms of variance in the mean-squared error across different values of spillover effects. This is suggestive that Louvain clustering with the largest number of clusters (Type 1) may be preferred in various applications.
5.1.3 Comparison between Bernoulli and Cluster design
To conclude, we present when a cluster design with our preferred clustering (Louvain Type 1) should be preferred over a Bernoulli design. Using Theorem, and the dense Messaging graph, 3.6 for Louvain clustering (Type 1), for any value of $\xi_{n}$ such that
$\xi_{n}^{1/2}\leq\bar{\xi}_{n}^{1/2},\quad\bar{\xi}_{n}^{1/2}=706,$
@@ 5.1.3 Comparison between Bernoulli and Cluster design
we should run a cluster experiment instead of a Bernoulli design. Because $\xi_{n}=\bar{\psi}/\bar{\phi}_{n}^{2}$ for $\lambda=1$, researchers should run a Cluster design if $\bar{\phi}_{n}^{2}\geq\bar{\psi}/\bar{\xi}_{n}.$ This implies that the researcher should run a cluster design if spillovers are larger than $\bar{\phi}_{n}\geq\bar{\psi}^{1/2}/706$. This comparison is suggestive that a cluster design with a Louvain clustering may be preferred over a Bernoulli design over a wide range of values of spillover effects (and outcomes variation $\bar{\psi}$). For binary outcomes, we should prefer cluster experiments for $\bar{\phi}_{n}\geq 0.003$.

### v3

@@ 4 Optimization over the choice of the clustering
Therefore, as a first step, we focus on minimizing $L_{c}(\xi)$ for each given $\xi$. Equation (9) solves a weighted min-cut problem with an additional penalization term that depends on the variance of the clusters’ size. The optimization problem in Equation (9) differs from the minimum normalized-cut problem (e.g., Shi and Malik, 2000; Ling and Strohmer, 2020), or related graph clustering problems (see Pokhilko et al., 2019; Parker et al., 2017; Zhang, 2023, and references therein) whose objective only depend on measure of cluster purity such as $\sum_{k=1}^{K}\sum_{i:c(i)=k}|j\in\mathcal{N}_{i}:c(i)\neq c(j)|/\sum_{i:c(i)=k}|\mathcal{N}_{i}|.$ Different from our proposal, standard min-cut problems use a different denominator (since the objective is not motivated by the bias of treatment effect estimators), and, most importantly, do not account for the additional variance component. Here instead, researchers should balance the bias and variance in the choice of the clusters. To the best of our knowledge, the problem in Equation (9) has not been studied before.
@@
Let $\mathbf{V}=\text{diag}(\mathbf{A}\mathbf{1}_{n})=\text{diag}(|\mathcal{N}_{1}(1)|,\ldots,|\mathcal{N}_{n}(1)|)$. Define the left-normalized Laplacian
@@
Let $\mathrm{tr}(\cdot)$ denote the trace operator. $c^{\star}$ minimizes (9) over the set of clusterings with exactly $K$ clusters if and only if
$c^{\star}\in\mathrm{arg}\max_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}\mathrm{tr}\Big{(}(n\mathbf{L}-\xi\mathbf{1}_{n}\mathbf{1}_{n}^{T})\mathbf{M}_{c}(K)\mathbf{M}_{c}^{T}(K)\Big{)}.$
@@
See Appendix D.8. ∎
Theorem 4.1 formalizes the optimization problem as a trace-optimization program for a fixed number of clusters $K$ and fixed $\xi$. Theorem 4.1 does not characterize a convex optimization program, but it provides a natural starting point to study convex relaxations of the proposed optimization problem. To obtain a convex relaxation, we relax the constraint on the matrix $\mathbf{X}(K)=\mathbf{M}_{c}(K)\mathbf{M}_{c}^{T}(K)$. We propose solving the following semi-definite programming (SDP) problem (for a given number of clusters $K$):
$\max_{\mathbf{X}(K)}\text{tr}(\mathbf{L}_{\xi}\mathbf{X}(K)),\quad\text{s.t.}\quad\text{diag}(\mathbf{X}(K))=\mathbf{1}_{n},\,\,\mathbf{X}(K)\succeq 0,\quad\mathbf{L}_{\xi;}=n\mathbf{L}-\xi\mathbf{1}_{n}\mathbf{1}_{n}^{T}.$ (11)
where Equation (11) defines a sequence of semi-definite optimization programs, each indexed by the number of clusters $K$ and multiplier $\xi$. The main distinction between Equation (9), and Equation (11) is that the matrix $\mathbf{X}(K)$ must be positive-definite instead of a matrix with binary entries. Let $\hat{\mathbf{X}}(K)$ be the solution of (11), for given $K$, that can be obtained using off-the-shelf optimization routines. We then apply the $K$-means algorithm on the first $K$ eigenvectors of $\hat{\mathbf{X}}(K)$ to retrieve the mapping $c$, whose properties are well studied in the literature (Von Luxburg, 2007). Finally, we compare solutions for different values of $K$ and choose the clustering with the largest objective.
@@
We can now describe the complete algorithm in Algorithm 1. We first posit the grid of values for $\xi$ that serves to approximate the Pareto frontier and a grid of values on the number of clusters $K_{n}$ we want to consider. A more fined grid provides a better approximation, at the cost of the optimization program scaling linearly in the size of each grid. For each value of $\xi$ and number of clusters $K_{n}$, we then solve a semi-definite program using off-the-shelf computational methods. For each value of $\xi$, we compute the best number of clusters $K_{n}$ by also minimizing $L_{c}(\xi)$ over the number of clusters. The resulting estimated clustering provides us with the set of clustering on a discretized Pareto frontier (approximated through the grid for $\xi$). Finally, we compute the mean-squared error optimal clustering by comparing the original objective in Equation (8) across all such estimated clustering over the grid of values for $\xi$ and $K_{n}$. The resulting optimization program is a simple sequence of semi-definite programs.
Algorithm 1 Causal Clustering
1:Adjacency matrix $\mathbf{A}$, $\underline{K}$, $\bar{K}$ smallest and largest number of clusters, expected relative magnitude of spillover effects $\bar{\phi}_{n}^{2}/\bar{\psi}$, weight on the variance $\lambda$, grid of values for $\xi\in\mathcal{S}\subset\mathbb{R}_{+}$.
2:for $\xi\in\mathcal{S}$ do
3:     for $K\in\{\underline{K},\cdots,\bar{K}\}$ do
4:a: Solve Equation (11) and obtain $\hat{\mathbf{X}}(K)$ as the minimized of Equation (11)
5:b: Retrive the clusters $c_{K}$ via $K$-means algorithm on the first $K$ eigenvectors of $\hat{\mathbf{X}}(K)$
6:c: Compute the objective function corresponding to the chosen clustering in (9)
7:     end for
8:     Save clustering with the lowest objective function in (9). Call this $\mathcal{C}(\xi)$.
9:end for
10:For each $\xi\in\mathcal{S}$ compute the objective in Equation (8) under each $\mathcal{C}(\xi)$.
11:Define $\mathcal{C}^{\star}\in\left\{\mathcal{C}(\xi),\xi\in\mathcal{S}\right\}$ the clustering with the lowest objective in (8). return $\mathcal{C}^{\star}$
Remark 8 (Spectral relaxation and comparison to spectral clustering).
@@
$\max_{\mathbf{U}}\text{tr}(\mathbf{L}_{\xi}\mathbf{U}(K)\mathbf{U}^{T}(K)),\quad\text{s.t.}\,\,\mathbf{U}^{T}(K)\mathbf{U}(K)=\mathbf{I}_{K},$
@@
$\text{tr}(\mathbf{L}_{\xi}\mathbf{U}(K)\mathbf{U}^{T}(K))=\text{tr}\left(\frac{\mathbf{L}_{\xi}+\mathbf{L}_{\xi}^{T}}{2}\mathbf{U}\mathbf{U}^{T}\right)$ (12)
The solution to the above problem $\hat{\mathbf{U}}(K)\in\mathbb{R}^{n\times K}$ is given by the matrix of top-$K$ eigenvectors of $(\mathbf{L}_{\xi}+\mathbf{L}_{\xi}^{T})/2$. Then we can perform a constrained $K$-means algorithm with equal-sized clusters to recover the clusters (Bradley et al., 2000). Finally, we iterate over the grid of values for $(\xi,K)$ to retrieve the optimal clustering.
This special case allows us to illustrate sharp distinctions with standard clustering methods such as spectral clustering (Von Luxburg, 2007). Once we fix the number of clusters $K$, the first distinction is which matrix we use to retrieve the clusters via K-means clustering. Spectral clustering would estimate the clustering through an eigendecomposition of the Laplacian matrix. Here instead, we use an eigendecomposition of the different matrix $\mathbf{U}\mathbf{U}^{\prime}$ that is a function of a left-normalized Laplacian and the matrix capturing the variance contribution of the estimated effect (implicit in $\mathbf{L}_{\xi}$). Instead of fixing the number of clusters, we then choose the clustering (and number of clusters) based on our different objective. ∎
Remark 9 (Reducing the computational complexity with a surrogate relaxation).
To reduce the complexity of Algorithm 1 it is possible to consider a single value of $\xi$ (instead of a grid of values as described in Algorithm 1) to minimize the mean-squared error. In particular, we can write $\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}+\frac{\lambda\bar{\phi}_{n}^{2}}{\bar{\psi}}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\Big{|}j\in\mathcal{N}_{i}:c(i)\neq c(j)\Big{|}\Big{)}^{2}\leq\frac{1}{\xi}L_{c}\left(\xi\right),$ for $\xi=\frac{\bar{\psi}}{\lambda\bar{\phi}_{n}^{2}}$, where $\frac{1}{\xi}L_{c}\left(\xi\right)$ is monotonically decreasing in $\xi$. Therefore, minimizing $\frac{1}{\xi}L_{c}(\xi)$ for any $\xi\leq\frac{\bar{\psi}}{\lambda\bar{\phi}_{n}^{2}}$ serves as a surrogate objective for the mean-squared error objective, where such surrogate imposes a larger weights to the bias relative to the variance. ∎
@@ 5 Empirical illustration and numerical studies
In this section, we illustrate the properties of the procedure in two empirical applications, one using unique data from the Facebook friendship and messaging graphs and one using data from an experiment conducted in rural China by Cai et al. (2015). We provide further evidence supporting our theoretical analysis using simulated networks in Appendix C.
5.1 Comparisons between existing clusterings at Faceboook
First, we focus on evaluating existing clusters currently implemented on Facebook at scale. [^17] We focus on existing clusterings (instead of implementing the causal clustering proposed here), as implementing new clustering methods requires changes to the entire engineering pipeline as described in Karrer et al. (2021), outside the scope of this paper. In Section 5.2 we turn to the implementation of Causal Clustering.
Bias and variance computations We consider two clustering algorithms implemented at scale on social networks owned by Meta: Louvain algorithm (Blondel et al., 2008), and Balanced Partitioning (Kabiljo et al., 2017). Louvain and Balanced Partitioning clustering produces a hierarchical clustering structure with a growing number of clusters. We consider three “types” of such algorithms, corresponding to three levels in the clustering structure hierarchy (i.e., different numbers of clusters), defined as Type 1, 2, and 3, respectively. For Louvain, higher-order types denote more clusters, whereas it is the opposite for Balanced Partitioning. For each clustering, we also report $\log(n/K_{n})$ in Table 3.
@@ 5.1 Comparisons between existing clusterings at Faceboook
We refer to the worst-case bias and variance weighted by $\xi$ as $b_{n}(\mathcal{C}_{n}),\frac{\xi}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}.$
@@
Results Table 3 in the appendix collects each of these algorithms’ worst-case bias and variance and the number of clusters per individual. A larger number of clusters increases the bias but decreases the variance throughout all graphs. (The bias increases for sparser networks because the denominator $|\mathcal{N}_{i}|$ decreases.) Louvain algorithm dominates Balanced Partitioning uniformly in variance and bias, suggesting that practitioners may prefer Louvain.
@@
We study different types of Louvain algorithms (each having a different number of clusters) in Figure 1. Figure 1 reports the worst-case mean squared error for different graphs, Louvain algorithms, and degrees of sparsity. For $\xi\approx 1$, the Louvain algorithm with the largest number of clusters (Type 3) dominates the other two algorithms in all but one case.
@@
As $\xi$ increases, Louvain clustering with the largest number of clusters (Type 3) is optimal for a dense graph but sub-optimal as the graph becomes more sparse. This result is intuitive: for denser networks, a larger number of clusters may best control the estimator’s bias relative to the bias induced by other clustering algorithms. However, we observe that Type 1 is the most robust in terms of variance in the mean-squared error across different values of spillover effects. This is suggestive that Louvain clustering with the largest number of clusters (Type 1) may be preferred in various applications.
@@
Comparison between Bernoulli and Cluster design To conclude, we study when a cluster design with our preferred clustering (Louvain Type 1) should be preferred over a Bernoulli design taking $\lambda=1$. Using Theorem 3.4, and the dense Messaging graph, Theorem 3.4 for Louvain clustering (Type 1), for $\sqrt{\frac{\bar{\psi}}{\bar{\phi}_{n}^{2}}}\leq\bar{\xi}^{1/2},\quad\bar{\xi}^{1/2}=706,$ we should run a cluster experiment instead of a Bernoulli design. Therefore, researchers should run a Cluster design if $\bar{\phi}_{n}^{2}\geq\bar{\psi}/\bar{\xi}.$ This implies that the researcher should run a cluster design if spillovers are larger than $\bar{\phi}_{n}\geq\bar{\psi}^{1/2}/706$. This comparison is suggestive that a cluster design with a Louvain clustering may be preferred over a Bernoulli design over a wide range of values of spillover effects (and outcomes variation $\bar{\psi}$). For binary outcomes, we should prefer cluster experiments for $\bar{\phi}_{n}\geq 0.003$.